Consider the following joint probability density function of the random variables X and Y :

3x−y , 1 < x < 3, 1 < y < 2, f(x, y) = 9 0, elsewhere.

(a) Find the marginal density functions of X and Y .

(b) Are X and Y independent?

(c) Find P(X > 2).
Answer #1

solution: Given that . (a) x 212 The marginal density function of agt and y fx(x)=sť vy) ay -$ (32-8) dy *{334-2), = yg [667-fr (x)= (6-Y) (14943) ( 1Y22) b) Note that fo (4) fy (Y)= 5(22-1)= (6-)=f(y) - » x and y are not independent. LC) PCX72) = $$